OpenStudy (anonymous):

]Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

iam sure

OpenStudy (anonymous):

y - 3 = -6(x - 2)

OpenStudy (anonymous):

y - 3 = -6(x - 2